I just need the answer for #5. I already calculated 1-4. Thank you!
Cornerstone Exercise 2.5 Costs of Services Jean and Tom Perritz own and manage Happy Home Helpers, Inc. (HHH), a house-cleaning service. Each cleaning (cleaning one house one time) takes a team of three house cleaners about 1.5 hours. On average, HHH completes about 15,000 costs are associated with the total cleanings: cleanings per year. The following total Direct materials Direct labor Variable overhead Fixed overhead S 27,000 472,500 15,000 18,000 Required: Calculate the prime cost per cleaning. 2. Calculate the conversion cost per cleaning. . Calculate the total variable cost per cleaning Calculate the total service cost per cleaning har ir rent on the office that Jean and Tom use to run HHH increased by $1,500? Explain the impact on the following: a. Prime cost per cleaning b. Conversion cost per cleaning c. Total variable cost per cleaning d. Total service cost per cleaning 5,
exercise z.c 2 2 I 5000
Answer #1

Answer

  • The main thing that you need to know that here Rent on the office is part of Fixed Overhead.
  • Requirement 5

--Part ‘a’
Prime Cost = Direct materials + Direct Labor.
Since Rent is part of Fixed Overhead, increase in rent will NO EFFECT on Prime cost per cleaning.

--Part ‘b’
Conversion cost = Direct labor + Overheads
= 472500 + 15000 + 18000 + 1500 increase = $ 507,000
New conversion cost per cleaning = $ 507000 / 15000 = $ 33.80
Hence, increase in rent will lead to INCREASE in Conversion cost per cleaning by $ 0.10 [33.80 – 33.70]

--Part ‘c’
Rent is a FIXED COST.
Hence, increase in Rent will not NO EFFECT on Total Variable cost per cleaning.

--Part ‘d’
Total Service cost = 27000 + 472500 + 15000 + 18000 + 1500 = $ 534,000
New Service cost per cleaning = 534000 / 15000 = $ 35.60
Hence, increase in Rent will lead to INCREASE in Total service cost per cleaning by $ 0.10 [35.6 – 35.5]